Five times the sum of the first and third of three consecutive even integers is ten more than the product of nine and the second integer. Find the third integer [i.e. the largest of the three].
Answer by stanbon(75887)

Five times the sum of the first and third of three consecutive even integers is ten more than the product of nine and the second integer. Find the third integer [i.e. the largest of the three].
---------------
1st: 2x
2nd: 2x+2
3rd: 2x+4
---------------
EQUATION:
5(2x + 2x+4) = 9(2x+2)+10
20x+20 = 18x+28
2x = 8
x = 4
----------
largest = 2x+4 = 12
=============
Cheers,
Stan H.